36X^2+30X-36 is the product of 6X-4 and what other factor?
1
Expert's answer
2012-08-07T09:20:41-0400
36X^2+30X-36=6(6x^2+5x-6)
6x^2+5x-6=0
x1,2=(-5+-sqrt(25+144))/12=(-5+-sqrt(169))/12=(-5+-13)/12
so x1=(-5+13)/12=8/12=2/3
x2=(-5-13)/12=-18/12=-3/2
Therefore 6x^2+5x-6=6(x-2/3)(x+3/2)=(3x-2)(2x+3)
So 36X^2+30X-36=6(6x^2+5x-6)=6(3x-2)(2x+3)=(6x-4)(6x+9)
Thus, the second factor is (6x+9)